OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−2, 4) and a directrix of y = 6. Put the equation in standard form. f(x) = one fourthx2 − x + 4 f(x) = −one fourthx2 − x + 4 f(x) = one fourthx2 − x + 5 f(x) = −one fourthx2 − x + 5

OpenStudy (anonymous):

Im getting B

OpenStudy (anonymous):

when focus and directrix are given use the deinition of the parabola or locus method.. let (x,y) be the point on the reqd parabola then distance of (x,y) from (-2,4) is sqrt ( (x+2)^2+(y-4)^2 ) and distance of (x,y) from y=6 is (y-6) units so by definition of parabola we have sqrt ( (x+2)^2+(y-4)^2 ) = (y-6) squaring both sides we have (x+2)^2+(y-4)^2 =(y-6)^2 and now just simplify

OpenStudy (anonymous):

so yes B

OpenStudy (anonymous):

yay :) Can you help with one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (0, -4) and a directrix of y = 4. (x - x)^2 + (y - 4)^2 = (x - 0)^2 + (y - -4)^2 y^2 -8y + 16 = x^2 + y^2 + 8y + 16 -16 y = x^2 y = - x^2 / 16 (or (-1/16)x^2

OpenStudy (anonymous):

so what do u think it is

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

hope i helped

OpenStudy (anonymous):

correct haha, thank you you are a lifesaver!!! :)

OpenStudy (anonymous):

your welcome